OpenStudy (anonymous):

A smaller commuter airline flies to three cities whose locations form the vertices of a right triangle. The totaly flight distance (from city A to city B to city C and back to city A) is 1400 miles. It is 600 miles between the two cities that are furthest apart. What are the other two distances between the cities?

OpenStudy (anonymous):

This requires use of the pythagorean theorem (a^2 + b^2 = c^2). I'll define the distance between cities A and B as a, the distance between cities B and C as b, and the distance between A and C as c. We find that c is the hypotenuse of this triangle if it has the greatest distance (600 miles). a^2 + b^2 = 600^2 a^2 + b^2 = 360000 We have an issue here. We don't have the distances a and b. To find that, we use the fact that a + b + c = 1400 miles given in the problem. Since c = 600, a + b = 800. Solve for b = 800 - a. Use this as b in the pythagorean theorem. a^2 + (800 - a)^2 = 360000 a^2 + (800 - a)(800 - a) = 360000 a^2 + 640000 - 1600a + a^2 = 360000 2a^2 - 1600a + 280000 = 0 a^2 - 800a + 140000 = 0 a^2 - 800a = -140000 a^2 - 800a + 160000 = 20000 (completing the square, we take -800/2 = -400 and square that to get 160000; then we add it to both sides) (a - 400)^2 = 20000 a - 400 = 100sqrt(2) a = 400 + 100sqrt(2) = 541 miles b = 800 - (400 + 100sqrt(2)) = 400 - 100sqrt(2) b = 259 miles

OpenStudy (anonymous):

well, whats the distance from city C to city A then? O.o

OpenStudy (anonymous):

That is the 600 miles given in the problem.

OpenStudy (anonymous):

But on the picture it says the 600 miles is from City A to City B O_O

OpenStudy (anonymous):

Oh, I'm sorry. I don't have a picture to look at, so I just chose the 600 miles to be from city C to city A. It can be from city A to city B if that's what the picture says. The 541 miles and 259 miles just represent the other two distances. You'll have to put them where they fit.

OpenStudy (anonymous):

Well that makes sense. Thank you so much!